Shift report 20030206 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Feb 06 2003 - 08:08:29 EST

  • Next message: Ian Bearden: "re:re-re-re-reduction of Au+Au data"
    ------------------------------------------------------------
    Date: Thu Feb  6 08:08:29 2003, Shift: 00:00-08:00
    Supervisor: ZM, with:  -
    Log book#14 pages 1 - 2
    
    Shift Summary:
    ------------------------------------------------------------
    1.We vote with PHOBOS to dump beam (not accetable high rate of TOFW. New store at 4:21 - good TOFW/ZDC ratio was obtained at 5:15.  Stil running.
    
    2.Run # 7809  520 K for trigger # 3 (MRS 40deg B1050).
      Run # 7810  new settings: MRS at 35 ded., B500
                                 FS at  8 deg., B1/15  
    ZDC=3.4K, TOF=42K and going down
    
    3. Cannot open mainframe window for ffs1440, bfs1440 and mrs1440 (despite BADTRIP was not running
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Thu Feb 06 2003 - 08:12:26 EST